OpenStudy (anonymous):

find theta, cos(theta)=5/3

OpenStudy (anonymous):

\[\theta = \cos^{-1} \frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

i tried that and it said error on my calc

OpenStudy (anonymous):

yep!!!! your angle can't be greater than 1. It's wrong question